OpenStudy (helpmeplease!!!):

what is 7+3(x+9)/2=15? Thanks

OpenStudy (helpmeplease!!!):

what is x? :)

OpenStudy (helpmeplease!!!):

and if you have some spare time could you also help me with 7(5x+8)-9=40

OpenStudy (anonymous):

7+3(x+9)/2=15 mulitply both sides by 2 7+3(x+9)=30 add 7&3 10(x+9)=30 divide both sides by 10 x+9=3 subrtact 9 from both sides x=-6

OpenStudy (anonymous):

7(5x+8)-9=40 add 9 on both sides 7(5x+8)=49 divide 7 on both sides 5x+8=7 subtract 8 on both sides 5x=-1 divide by 5 x=-1/5

OpenStudy (anonymous):

@heatherfeather Your answer is incorrect. It's isn't -6. You are supposed to distributive 3 to (x + 9)

OpenStudy (anonymous):

@micahwood50 ah okay. so it would be: 7+3(x+9)=15 muliply both sides by 2 7+3(x+9)=30 distribute the 3 to the x&9 7+3x+9=30 add the 7&9 16+3x=30 subtract 16 from both sides 3x=14 divide by 3 x= 4.7

OpenStudy (helpmeplease!!!):

Thank you heatherfeather for helping me and thank you micahwood50 for helping heatherfeather. :D You guys were a great help.